# Collection: UAV-Based Wireless Multi-modal Measurements from AERPAW Autonomous Data Mule (AADM) Challenge in Digital Twin and Real-World Environments

### v1

@@ Abstract
In the first stage (DT evaluation), each team’s proposed solution was run in the DT environment with three separate UAV missions. The overall competition workflow, including the DT and RW testbed stages, is illustrated in Fig. 2. Each mission corresponds to different data-volume configurations across AERPAW BSs to make the challenge more complex and to optimize the trajectory with heterogeneous traffic demands. Four data-volume configurations, denoted as D1, D2, D3, and D4, are defined across four BSs BS1, BS2, BS3, and BS4, to introduce heterogeneous traffic demands. The first flight uses the configuration (600, 600, 600, 200) Mbits, the second flight uses (100, 400, 400, 100) Mbits, and the third flight uses (20, 50, 400, 30) Mbits assigned to BS1 through BS4, respectively. In each case, the UAV aims to maximize the total downloaded data within 500 seconds. In total, 45 flights (15 teams times 3 flights) are run in stage 1.
@@
Finally, the dataset can be used to explore the scalability of data-mule operations in next-generation aerial networks. Researchers can simulate scenarios involving multiple UAVs collecting data in parallel under different BS data-demand configurations and validate their results with the DT and RW testbed data logs.
@@
All source code and post-processing scripts employed in the cleaning, merging, and analysis of the dataset are made publicly available alongside the dataset at Dryad (DOI: 10.5061/dryad.7d7wm3898). The dataset includes MATLAB source code to merge log files to generate datasets in CSV format, compute SNR and cumulative download statistics, and recreate the trajectory, altitude, speed, and power versus distance plots. Other scripts are employed to process the LoRa gateway logs to obtain RSSI and SNR statistics, convert the Fortem radar JSON data to a MATLAB format, and combine the Fortem, Keysight RF sensor data, and UAV ground truth data while also computing localization errors. Other LTE processing scripts are included to generate the RSRP versus distance plots from the raw IQ recordings.
@@
The authors thank NI, Keysight Technologies, AUVSI NC, and 6GNC for sponsoring the challenge awards. The authors have no conflicts of interest.
    M. Hossen and C. Dickerson led the manuscript preparation and data analysis. I. Guvenc guided the revision of the paper. All other authors contributed through manuscript review. Eleven independent participating teams contributed to algorithm development. These teams were led by the following representatives: A. M. Alam, W. Bjorndahl, F. Solis, S. Javed, A. Kamath, X. Tang, J. J. Sadique, K. L. Hermstein, K. A. Mahmud, J. A. S. Viloria, E. Tucker, S. Hawkins, Y. Cui, A. Dey, Y. Liu, A. Gurbuz, J. Camp, R. Ahmad, J. van der Merwe, A. I. Mohamed, G. Zussman, M. Kurum, N. Kamesh, Z. Guan, D. Pados, and G. Skilvanitis. The AERPAW platform development and the execution of the field experiments were carried out by O. Ozdemir, A. Gurses, T. Zajkowski, M.R. Sarbudeen, I. Guvenc, M. Sichitiu, M. Mushi, and R. Dutta.

### v2

@@ Abstract
In the first stage (DT evaluation), each team’s proposed solution was run in the DT environment with three separate UAV missions. The overall competition workflow, including the DT and RW testbed stages, is illustrated in Fig. 2. Each mission corresponds to different data-volume configurations across AERPAW BSs to make the challenge more complex and to optimize the trajectory with heterogeneous traffic demands. Four data-volume configurations, denoted as D1, D2, D3, and D4, are defined across four BSs: BS1, BS2, BS3, and BS4, to introduce heterogeneous traffic demands. The first flight uses the configuration (600, 600, 600, 200) Mbits, the second uses (100, 400, 400, 100) Mbits, and the third flight uses (20, 50, 400, 30) Mbits assigned to BS1 through BS4, respectively. In each case, the UAV aims to maximize the total downloaded data within 500 seconds. In total, 45 flights (15 teams times 3 flights) are run in stage 1.
@@
It can be used to explore the scalability of data-mule operations in next-generation aerial networks. Researchers can simulate scenarios with multiple UAVs collecting data in parallel under different BS data-demand configurations and validate their results with the DT and RW testbed data logs.
@@
All source code and post-processing scripts employed in the cleaning, merging, and analysis of the dataset are made publicly available alongside the dataset at Dryad (DOI: 10.5061/dryad.7d7wm3898). The dataset includes MATLAB source code to merge log files to generate datasets in CSV format, compute SNR and cumulative download statistics, and recreate the trajectory, altitude, speed, and power-versus-distance plots. Other scripts are used to process the LoRa gateway logs to obtain RSSI and SNR statistics, convert the Fortem radar JSON data to a MATLAB format, and combine the Fortem, Keysight RF sensor data, and UAV ground truth data while also computing localization errors. Other LTE processing scripts are included to generate the RSRP versus distance plots from the raw IQ recordings.
@@
Special thanks to NI, Keysight Technologies, AUVSI NC, and 6GNC for sponsoring the challenge awards.
This work was supported by the NSF under Grant CNS-1939334 and in part by CNS-2450593 (AERPAW). Additional support for participating teams was provided in part by NSF Grants 2332661, OAC-2512931, 2346555, CNS-2144297, CNS-2148128, CNS-2332662, CNS-2117822, EEC-2133516, CNS-2440756, SWIFT-2229563, DGE-2137100, and by HEC, PK.
M. S. Hossen prototyped the AADM challenge software framework, led the dataset curation process, processed and analyzed the USRP measurements, and prepared the initial manuscript draft. C. Dickerson processed and validated the radar and RFS datasets and assisted with manuscript preparation. O. Ozdemir processed and analyzed the LoRa and LTE measurements and validated the dataset.
The AERPAW platform development and the execution of the field experiments were carried out by O. Ozdemir, A. Gurses, M. R. Sarbudeen, T. Zajkowski, I. Guvenc, M. Sichitiu, M. Mushi, and R. Dutta.
All remaining authors participated in the AADM challenge and contributed to algorithm development in the digital twin environment, experimental execution, data validation, and critical review of the manuscript.
The authors declare no conflicts of interest.